OpenStudy (anonymous):

evaluate 1og base 7 343 can someone please explain how i evaluate this?

OpenStudy (anonymous):

343 = 7^3...help?

OpenStudy (anonymous):

log7 343\[\log_{7} 343\]

OpenStudy (anonymous):

If u have log_a a^x the answer is x So log_7 7^3 = 3

OpenStudy (anonymous):

I'd evaluate it with a calculator. But that basically means the number you raise 7 to to get 343.

OpenStudy (anonymous):

Get it?

OpenStudy (anonymous):

i dont know how to put it into my caculator

OpenStudy (anonymous):

log_{7} 343 = x means 7^x=343

OpenStudy (anonymous):

eg log_10 100 = log_10 10^2 = 2

OpenStudy (anonymous):

It is the reverse of the other questions u were doing...

OpenStudy (anonymous):

7 to the what power = 343

OpenStudy (anonymous):

\[\log7^{343}\]?

OpenStudy (anonymous):

no the 7 is the base

OpenStudy (anonymous):

Quiz

OpenStudy (anonymous):

Calculators only have log base 10 and natural log, which is log base e. So you have to use a sweet math trick to get log base 7. If you want log_7 343, you enter (log_10 343) /log_10 7

OpenStudy (anonymous):

2 way you can solve

OpenStudy (anonymous):

Or in general, log_x y can be entered as log_10 y / log_10 x

OpenStudy (anonymous):

\[\frac{\log 343}{\log 7}=3\]

OpenStudy (anonymous):

\[x = \log_ay\] means (is the same as saying) \[y=a^x\]

OpenStudy (anonymous):

or: \[\log_7 ^{x}= 342\]\[7^x=343\]\[7^x=7^3\] x=3\[\log_7^{({3})}=343\]

OpenStudy (anonymous):

woulndt that mean the answer is 49?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

7^3=343